Please do not answer if you are not an expert. You will not like the review if you guess.

Please explain with detailMatch the surfactants to the HLB you would expect for them, using the expression: HLB 7+(hydrophilic group numbers) - 2(hydro
Answer #1

1) HO(CH2CH2O)10CH2OH

HLB = 7 + \Sigma (hydrophilic group numbers) - \Sigma (hydrophobic group numbers)

sum of Hydrophilic group numbers = 1.9 (OH) + 0.33X 10 (CH2CH2O) + 1.9 (OH) = 7.1

Sum of hydrophobic group number = 0.475

HLB = 7 + (7.1) - (0.475) = 13.625 = 13.6

2) CH3(CH2)13COOK

HLB = 7 + \Sigma (hydrophilic group numbers) - \Sigma (hydrophobic group numbers)

sum of Hydrophilic group numbers = 19.1 (-COOK) = 19.1

Sum of hydrophobic group number = 0.475 (CH3) + 13 X 0.475 (CH2) = 6.65

HLB = 7 + (19.1) - (6.65) = 19.45 = 19

3) N(CH2CH2CH2OH)3

HLB = 7 + \Sigma (hydrophilic group numbers) - \Sigma (hydrophobic group numbers)

sum of Hydrophilic group numbers = 9.4 (-N+3) + 3 X 1.9 (-OH) = 15.1

Sum of hydrophobic group number = 9 X 0.475 (CH2) = 4.275

HLB = 7 + (15.1) - (4.275) = 17.825 = 17.8

4) CH3(CH2)17OH

HLB = 7 + \Sigma (hydrophilic group numbers) - \Sigma (hydrophobic group numbers)

sum of Hydrophilic group numbers = 1.9 (-OH) = 1.9

Sum of hydrophobic group number = 0.475 (CH3) + 17 X 0.475 (CH2) = 8.55

HLB = 7 + (1.9) - (8.55) = 0.35 = 0.4